OpenStudy (anonymous):

Which expressions are equivalent to (5^3x5^2)^4 a.25^20 b.(25^5)4 c.none of the above if possible can you explain please.

rishavraj (rishavraj):

\[a^x \times a^y = a^{x+y}\] and also \[(a^x)^y = a^{xy}\]

OpenStudy (anonymous):

huhh?

rishavraj (rishavraj):

hmmm \[5^x \times 5^y = 5^{x+y}\]

OpenStudy (anonymous):

oh, so none of the above

rishavraj (rishavraj):

hold on...think over it and also I told u \[(a^x)^y = a^{xy}\] so whts the value of \[5^3 \times 5^2 \]

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

are you showing mw how to get the answer, because im not understadning

rishavraj (rishavraj):

I am making u understand -_- as I told u \[a^x \times a^y = a^{x+y}\]

OpenStudy (anonymous):

ok sorry, I get that now.